There is a number written on the board. One student increased the number by 23, but another student decreased the number by 1. The first student’s result was 7 times greater than the result of the second student. What is the number written on the board?
Answer by ikleyn(52775)

.
There is a number written on the board. One student increased the number by 23, but another student decreased the number by 1.
The first student’s result was 7 times greater than the result of the second student. What is the number written on the board?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

x + 23 = 7(x-1).


x + 23 = 7x - 7,


6x = 23+7  --->  6x = 30  --->  x = 5.